40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Sep 12 20:19:25 2016, Lydia, Update, General, MC REFL beam splitter not replaced 
    Reply  Wed Sep 14 19:41:23 2016, Johannes, Update, General, PSL back to high power 
Message ID: 12485     Entry time: Mon Sep 12 20:19:25 2016     Reply to this: 12493
Author: Lydia 
Type: Update 
Category: General 
Subject: MC REFL beam splitter not replaced 

The beam splitter that directs light into the MC REFL photodiode has not been replaced; there is still a mirror there. Gautam suggested we wait to replace it until the PSL shutter is open so the beam can be aligned. However, this must be done before going to high power.

GV addendum: What I suggested was to try and recover the arm alignment using the current low power configuration after pumpdown - since we were well aligned just before pumpdown, we should be able to recover this alignment pretty easily at low power. After locking both arms and running the dither align (also center all Oplevs), we can go ahead do the following:

  • Replace mirror in MC Refl path with 10% reflection BS (Johannes, Lydia and I confirmed that this is on the AP table earlier today). Then align the reflected beam onto the PD using the tiny mirror
  • Replace HR mirror in Transmon path at the EY table
  • Replace ND filters on Transmon QPDs at EX and EY tables
  • Repalce ND filter on Transmon CCD at EY table
  • Revert MC autolocker to the nominal version instead of the low power version we have been using during the vent
  • Turn up MC to nominal power by rotating the wave plate on the PSL table - confirm that we have nominal levels by measuring with power meter
  • Recover single arm locks, green beatnotes etc at nominal operating conditions
     
ELOG V3.1.3-